Is the quadratic form x12+x22+2x1x3+4x1x3+3x32 quadratic? Justify your answer


1
Expert's answer
2021-01-11T17:35:05-0500

A quadratic form is a polynomial with terms all of degree two. 

Quadratic forms are homogeneous quadratic polynomials in "n" variables.

In our case "n=3," and we have ternary quadratic form

"f(x_1, x_2, x_3)=1\cdot x_1^2+0\cdot x_1x_2+6\cdot x_1x_3+1\cdot x_2^2"

"+0\cdot x_2x_3+1\cdot x_3^2"

The quadratic form "f(x_1, x_2, x_3)" is determined by the matrix

"A=\begin{bmatrix}
 1 & 0 & 3 \\
 0 & 1 & 0 \\
 3 & 0 & 1
\end{bmatrix}"